OpenStudy (gigi1103):

What is the value of the function y = 2x + 3 ⁢when x=−1 ? check my answer?

OpenStudy (gigi1103):

is it positive one?

OpenStudy (redheadangel):

1. Solve what's inside the paranthesis 2. multiply(*) / divide (:) 3. Add(+) / subtract (-) y = 2x - 3 = 2(-1) - 3 = -2 - 3 = -5

OpenStudy (redheadangel):

whoops i thought it said -3

OpenStudy (redheadangel):

disregard that

OpenStudy (redheadangel):

it's positive 1

OpenStudy (gigi1103):

Thank you thank you

OpenStudy (redheadangel):

of course!